SITE 1043

(Proposed Site CR-6)

Hole 1043A
Hole 1043A was spudded at 1545 hr on 8 December with the bit at 4322.5 mbrf. Core 170 1043A-1H recovered 8.01 m of sediment. APC Core 170-1043A-2H failed to bleed off pressure, indicating incomplete stroke in the firm seafloor material. XCB coring began with core 3X and continued through Core 170-1043A-18X at a depth of 167.1 mbsf. A small amount of fill and slightly elevated pump pressure was identified at 157.5 mbsf, after Core 170-1043A-17X. A single 30 bbl sweep of Sepiolite drilling mud cleaned up the hole and no further problems were encountered. A rapid increase in ROP occurred while cutting Core 170-1043A-18X, which only took five minutes to cut. Although the formation softened, the material remained quite stiff. The APC barrel was deployed for Core 170-1043A-19H (167.1 to 176.6 mbsf) in an attempt to get an undisturbed core sample. The barrel did achieve full stroke and recovered 9.83 m of high quality core. Because the sediment was even more highly compacted than on the previous core, and it was deemed unwise to advance the hole below the décollement with the core liner in the pipe, coring was resumed with the XCB system. Coring continued with the XCB to a total depth of 282.3 mbsf. Coring was abandoned at this point to allow time for two final LWD holes at Sites 1043 and 1042. The hole was displaced with 75 bbls of 10.5 lb/gal mud. As the pipe was tripped, the vessel was offset 50 m to the southeast for spudding Hole 1043B. The bit arrived at the rig floor by 0545 hr 11 December 1996, ending Hole 1043A.

A single temperature measurement with the DVTP was attempted after Core 170-1043A-4X at 33.4 mbsf. However, the data was bad (formation too hard). Subsequent temperature measurements were canceled.

Evidence of the presence of hydrate nodules was seen beginning with Core 170-1043A-6X. Hydrocarbon levels were highly variable and dropped significantly below the top of the underthrust zone as identified in Core 170-1043A-17X at ~150 mbsf. Methane values measured by headspace analysis ranged from 81 to 237,615 ppm above the décollement and ranged from 6 to 165 ppm below it. Ethane values ranged from 2 to 48 ppm above the décollement and nothing below Core 170-1043A-18X (167.1 mbsf). Propane measured 0-6 ppm and no higher hydrocarbons were identified.

Hole 1043B
Hole 1043B was to be the first of two back-to-back LWD holes. The jars were pressure tested (low 500 psi and high 3000 psi) prior to making up the LWD BHA with a tricone drill bit. The CDR and CDN LWD tools were made up as part of a standard BHA including the McCullough mechanical drilling jars. The drill string was tripped to the seafloor, filling the pipe every 40 stands. Hole 1043B was spudded at 1645 hr 11 December 1996. Drilling continued at a target rate of 25 m/hr until 0430 hr 12 December 1996 when the driller was advised by Anadrill/LDEO that 35 m/hr was acceptable. At a depth of 196.0 mbsf, the target ROP was increased accordingly. Drilling continued without incident to a total depth of 4803.3 m (482.3 mbsf), when a drastic reduction in ROP as well as an increase in torque indicated that the target reflector had been reached. Basement for the site was projected at ~490.0 +10 mbsf. It was assumed that the hole stopped on basement or indurated sediments directly overlying basement.

No problems were experienced penetrating the décollement and only a single 30 bbls Sepiolite mud sweep was required during the drilling. The pill was circulated at a depth of 453.4 mbsf where a slight increase in pump pressure was recognized by the driller. The hole was displaced with 140 bbls of 10.5 lb/gal mud before being abandoned. The vessel was then offset back to Site 1042. The bit cleared the rotary table at 0345 hr 13 December 1996, thereby ending Hole 1043B.

To 170 Operations Chronology

To 170 Table of Contents

Publications Homepage

ODP Homepage